OpenStudy (anonymous):

Describe the transformation from f(x)=a^x to f(x)=b^x where 1

OpenStudy (dumbcow):

graphs have same shape but slope is greater with the a^x (it increases at faster rate)

OpenStudy (anonymous):

I was asking for more along the lines of vertical stretch, horizontal stretch, etc.

OpenStudy (perl):

its a horizontal stretch, relatively to the parent. for example y = 2^x , y = .5 ^ x

OpenStudy (anonymous):

by what factor?

OpenStudy (perl):

a ^[ x * ln b / ln a ] = b^x

OpenStudy (anonymous):

how did you get that?

OpenStudy (perl):

look at easier examples, comparing y = 2^x and y = 2^(2x)

OpenStudy (perl):

ok here is a general rule

OpenStudy (perl):

you ready ?

OpenStudy (perl):

if you start with f(x) , f (ax) does a horizontal stretch/shrink by a factor of 1/a. it depends on a, and we assume a > 0 . if 0<a<1, then its a horizontal stretch because if a < 1 then 1/a > 1 , and if a > 1 , then it is a horizontal shrink because 1/a < 1

OpenStudy (perl):

so it becomes clear if its a stretch or a shrink because of the factor itself. if the factor is less than 1, then its a shrink. if its greater than 1 its a stretch

OpenStudy (perl):

now we started with f(x) = a^x , then f ( ln b/ ln a * x ) = a ^ (ln b / ln a * x ) = b^x . so f(x) is stretched or shrinked by a factor of ln a / ln b (the reciprocal )

OpenStudy (perl):

the important thing is that it is horizontal

OpenStudy (perl):

you there ?

OpenStudy (anonymous):

I understand the picture but I don't understand how that translated into lnb/lna

OpenStudy (anonymous):

the 1st picture

OpenStudy (perl):